# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v1

@@
Abstract.
Wi-Fi sensing has emerged as a transformative technology that leverages ubiquitous wireless signals to enable a variety of applications ranging from activity and gesture recognition to indoor localization and health monitoring. However, the inherent dependency of Wi-Fi signals on environmental conditions introduces significant generalization challenges—variations in surroundings, human positions, and orientations often lead to inconsistent signal features, impeding robust action recognition.
In this survey, we review over 200 studies on Wi-Fi sensing generalization, categorizing them along the entire sensing pipeline: device deployment, signal processing, feature learning, and model deployment. We systematically analyze state-of-the-art techniques, which are employed to mitigate the adverse effects of environmental variability. Moreover, we provide a comprehensive overview of open-source datasets such as Widar3.0, XRF55, and XRFv2, highlighting their unique characteristics and applicability for multimodal fusion and cross-modal tasks. Finally, we discuss emerging research directions—including multimodal approaches and the integration of large language models—to inspire future advancements in this rapidly evolving field. Our survey aims to serve as a valuable resource for researchers, offering insights into current methodologies, available datasets, and promising avenues for further investigation.
@@ Abstract.
copyright: acmlicensedjournal: CSURjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviews
1. Introduction
@@
Wi-Fi sensing has emerged as a transformative technology that exploits the pervasive Wi-Fi infrastructure for purposes beyond traditional communication. By examining fluctuations in Wi-Fi signals, particularly Channel State Information (CSI) and Received Signal Strength Indicator (RSSI), researchers have explored its potential across diverse applications, including activity recognition (Jiang et al., 2018b; Li et al., 2021d; Wang et al., 2017a, 2015; Tan et al., 2019; Zhu et al., 2017b; Wang et al., 2014; Chen et al., 2018; Adib and Katabi, 2013; Qian et al., 2017b), gesture recognition (Li et al., 2016b; Pu et al., 2013; Tan and Yang, 2016; Li et al., 2016c; Ali et al., 2015; He et al., 2015; Venkatnarayan et al., 2018; Abdelnasser et al., 2015; Ma et al., 2018; Wang et al., 2019a; Sun et al., 2015), indoor localization (Li et al., 2016a; Xie et al., 2015; Wang et al., 2016a; Vasisht et al., 2016; Qian et al., 2017a; Gjengset et al., 2014; Li et al., 2017; Xie et al., 2019; Qian et al., 2018; Kotaru et al., 2015), user authentication and identification (Korany et al., 2019; Shahzad and Zhang, 2018; Wang et al., 2019c, b; Shi et al., 2017; Zeng et al., 2016; Zhang et al., 2020c; Wang et al., 2016b), health monitoring (Zeng et al., 2020; Wang et al., 2017b; Zeng et al., 2018; Wang et al., 2016d; Zheng et al., 2016; Liu et al., 2014; Zeng et al., 2019), fallen detection (Wang et al., 2016f; Palipana et al., 2018; Wang et al., 2016c; Ding and Wang, 2020; Hu et al., 2021b), pose estimation (Jiang et al., 2020; Wang et al., 2019e, f; Ren et al., 2021, 2022; Zhou et al., 2023; Wang et al., 2024c, 2022b; Qian et al., 2023; Yan et al., 2024b), crowd counting (Zou et al., 2017, 2018b; Choi et al., 2022; Jiang et al., 2023; Xi et al., 2014), and more. Numerous survey papers have comprehensively summarized this flourishing research area, examining existing studies from various perspectives such as methodologies, application domains, and technical challenges (Tan et al., 2022; Hernandez and Bulut, 2022; Zhang et al., 2020a; Ma et al., 2019; Liu et al., 2019; Xiao et al., 2016; Li et al., 2021a; Jiang et al., 2018a; Yang et al., 2013; Yousefi et al., 2017).
Wi-Fi sensing systems face significant deployment challenges in real-world environments. This is primarily because Wi-Fi signals are intricately linked to their surroundings, so any alteration in the environment or the people within it can modify the signal characteristics. Even when the same person performs an identical action in the same setting, variations in position or orientation can result in inconsistent signal features, complicating reliable action recognition. To enable more widespread application of Wi-Fi sensing, researchers increasingly use advanced techniques like domain adaptation, domain generalization, and meta-learning to overcome these issues.
Recent studies have increasingly focused on enhancing the generalizability of Wi-Fi sensing techniques. For instance, Chen et al (Chen et al., 2023) adopts a cross-domain perspective, reviewing and categorizing over 60 papers that address challenges in Wi-Fi sensing. This work organizes various techniques—including domain-invariant feature extraction, transfer learning, and few-shot learning—that contribute to cross-domain robustness. In contrast, Wang et al (Wang et al., 2024a) zeroes in on the few-shot learning paradigm, providing a detailed overview of approaches such as transfer learning, metric learning, and meta-learning. Together, these studies highlight the significant progress made in applying advanced learning strategies to improve the adaptability and performance of Wi-Fi sensing systems.
Compared with previous works, this survey encompasses a significantly broader range of Wi-Fi generalization studies—over 200 papers in total. We have organized these studies along the entire Wi-Fi sensing pipeline: from device deployment and signal processing to feature learning and model deployment. This structured approach enables researchers to quickly locate the techniques that best address their specific challenges. Furthermore, we provide a detailed overview of open-source datasets, helping researchers efficiently identify the data needed to validate their algorithms. Finally, we offer a comprehensive analysis of future directions, including trends in multimodal approaches and large language models, to inspire new research ideas.
In summary, this work serves as a valuable resource for researchers seeking convenient access to methodologies, datasets, and innovative research strategies.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6em,align=left,fit=band,font=,, where level=2text width=8.2em,align=left,fit=band,font=,, where level=3text width=6.8em,font=,, where level=4text width=5em, font=,, [ Wi-Fi Human Sensing Generalization Categories, ver [ Hardware
@@
2.1) , fill=orange!5 [ Distributed Antennas
2.1.1), fill=orange!10 [ WiNDR (Qin et al., 2024), WiCross (Qin et al., 2023), WiSDAR (Wang et al., 2018) , leaf,fill=orange!10, text width=40em ] ] [ Distributed Devices
2.1.2),fill=orange!10 [ ReWiS (Bahadori et al., 2022), Widar 3.0 (Zheng et al., 2019), OneFi (Xiao et al., 2021), WiTraj (Wu et al., 2021b) , leaf,fill=orange!10, text width=40em ] ] [ Scaling Up Dataset
2.1.3),fill=orange!10 [ XRF55 (Wang et al., 2024b), Widar 3.0 (Zheng et al., 2019), Zhang et al.(Zhang et al., 2023h), MM-Fi(Yang et al., 2024a) , leaf,fill=orange!10, text width=40em ] ] ] [ Signal
@@
2.2), fill=violet!5 [ Signal Analysis
Methods (§2.2.1), fill=violet!10 [ WiHand (Lu et al., 2019), CLAR (Zhou et al., 2022a), Gait-Enhance(Yang et al., 2023a), Zhuo et al.(Zhuo et al., 2023), Abuhoureyah et al. (Abuhoureyah et al., 2024), leaf, fill=violet!15, text width=40em ] ] [ Advanced Handcrafted
Indicators (§2.2.2), fill=violet!10 [ WiHGR (Meng et al., 2021), PhaseAnti (Huang et al., 2021), DPSense (Gao et al., 2022a), HandGest (Zhang et al., 2022d), Yu et.al. (Yu et al., 2023), AF-ACT (Zhang et al., 2022e), RoSeFi (Peng et al., 2023),
Wang et.al. (Wang et al., 2024e), WiGesFree(Ding et al., 2024), LiKey (Peng et al., 2024), Bu et.al. (Bu et al., 2018), Zhang et.al. (Zhang et al., 2023e) , leaf, fill=violet!15, text width=40em ] ] [ Motion Indicators
2.2.3), fill=violet!10 [ Sharp (Meneghello et al., 2022), OneSense (Zhao et al., 2024d), Yin et al.(Yin et al., 2021), Wi-learner (Feng et al., 2022), Niu et al. (Niu et al., 2021), and LAGER (Chen et al., 2024a), WiLife (Li et al., [n. d.]),
Mini-Batch Alignment (van Berlo et al., 2023, 2022), WiTraj (Wu et al., 2021b), Widar 3.0 (Zheng et al., 2019). Zhang et al. (Zhang et al., 2023i), WiTransformer (Yang et al., 2023c),
Shi et al. (Shi et al., 2023), Bulugu (Bulugu, 2023). GaitID (Zhang et al., 2020d), OneFi (Xiao et al., 2021), XFall (Chi et al., 2024), WiGesture (Gao et al., 2021), Shi et al. (Shi et al., 2020a),
Wang et al. (Wang et al., 2022c), Widir (Wu et al., 2016), Wang et al. (Wang et al., 2016e), WiNDR(Qin et al., 2024), WiDIGR (Zhang et al., 2019a), Wi-PIGR (Zhang et al., 2021d),
WiDetect (Zhang et al., 2019b), Zhu et al. (Zhu et al., 2024), leaf, fill=violet!15, text width=40em ] ] [ Angle Indicators
2.2.4), fill=violet!10 [ Li et al. (Li et al., 2020), EasyTrack (Han et al., 2023), 3D-ID (Ren and Yang, 2022), leaf, fill=violet!15, text width=40em ] ] ] [ Feature
@@
2.3),fill=yellow!5 [ Domain Alignment
@@
2.3.1),fill=yellow!10 [ JADA (Zou et al., 2018a), CrossGR (Li et al., 2021b), WiHARAN (Wang et al., 2021a), i-Sample (Zhou et al., 2024), CsiGAN (Xiao et al., 2019), CycleGAN (Zhu et al., 2017a), Wi-Adaptor (Zhang et al., 2021e),
GESFI (Zhang et al., 2024a), Khattak (Khattak and Khan, 2022), DINN (Zhou et al., 2021), INDG-Fi (Yang et al., 2024b), freeGait (Yan et al., 2024c), CATFSID (Wei et al., 2024), ADA (Zinys et al., 2021), DAFI (Li et al., 2021c),
Shi et al. (Shi et al., 2020a), DATTA (Strohmayer et al., 2024), WiSR (Liu et al., 2023a), WiCAR (Wang et al., 2019d), Berlo et al. (Berlo et al., 2023), WiCross (Qin et al., 2023), alignment.Li et al. (Li et al., 2021e),
Hassan et al. (Hassan et al., 2024), WiSDA (Jiao et al., 2024), LAGER (Chen et al., 2024a), WiCAU (Cui et al., 2024), CrossFi (Zhao et al., 2024a), P3ID (He et al., 2024), AdapLoc (Zhou et al., 2020), WiAi-ID (Liang et al., 2023),
Chen et al. (Mehryar, 2023), Zhang et al. (Zhang et al., 2021b), Mini-Batch Alignment (van Berlo et al., 2023, 2022), MetaGanFi (Zhang et al., 2022b), CsiGAN (Xiao et al., 2019), cyclegan (Zhu et al., 2017a),
SIDA (Zhang et al., 2023b), Zhang et al. (Zhang et al., 2022a), EI (Jiang et al., 2018c), AirFi (Wang et al., 2022d)CrossID (Wu et al., 2021a), Wang et al. (Wang et al., 2024d), Wi-Piga (Hao et al., 2021), Wi-Dist (Zhang et al., 2024d),
UH-Sense (Chen et al., 2024b), Person-in-WiFi (Wang et al., 2019f), Elujide et al. (Elujide et al., 2022), SCTRFL (Zheng et al., 2020), PIAS (Xiao et al., 2024b), WiLISensing (Ding et al., 2020b) , leaf,fill=yellow!15, text width=40em ] ] [ Meta-learning
2.3.3),fill=yellow!10 [ MAML (Finn et al., 2017), Metaloc (Gao et al., 2023), RoMF (Zhang et al., 2024b), WiLiMetaSensing (Ding et al., 2021a), MetaGANFi (Zhang et al., 2022b), WiOpen (Zhang et al., 2024c), CSI-GDAM (Zhang et al., 2021a),
FewCS (Yang et al., 2023b), WiGesID (Zhang et al., 2022c), MatNet-eCSI (Shi et al., 2020b), LT-WiOB (Zhou et al., 2022c), TransferSense (Bu et al., 2022), DAFI  (Li et al., 2021c), MetaGANFi (Zhang et al., 2022b),
CLAR-Xiao et al. (Xiao et al., 2024a), DualConFi (Xu et al., 2022), SFTSeg  (Xiao et al., 2022), HAR-MN-EF (Shi et al., 2020c), DA-HAR (Sheng et al., 2023), Shi et al. (Wang et al., 2022e), WiDeo(Islam and Nirjon, 2021)
RF-Net (Ding et al., 2020a), Wi-learner (Feng et al., 2022), Reptile (Nichol, 2018), ML-WiGR (Gao et al., 2022b), TOSS (Zhou et al., 2022b), Zhang et al. (Zhang et al., 2020b), Huang et al. (Huang et al., 2022),
AFSL-HAR (Wang et al., 2021b), Wi-AM (Zhang et al., 2023j), Prototypical Network (Snell et al., 2017), MetaFormer (Sheng et al., 2024) , leaf,fill=yellow!15, text width=40em ] ] [ Embedding Learning
2.3.4),fill=yellow!10 [ FewCS (Yang et al., 2023b), ReWiS (Bahadori et al., 2022), WiGesID  (Zhang et al., 2022c), MatNet-eCSI (Shi et al., 2020b), LT-WiOB (Zhou et al., 2022c), TransferSense  (Bu et al., 2022), 3D-ID  (Ren and Yang, 2022),
WiDeo (Islam and Nirjon, 2021), CrossFi  (Zhao et al., 2024a), SFTSeg  (Xiao et al., 2022), CSI-GDAM (Zhang et al., 2021a), UniFi (Liu et al., 2024b), MaP-SGAN (Liang et al., 2024), Wi-Cro (Mao et al., 2024), CATFSID (Wei et al., 2024),
WiOpen (Zhang et al., 2024c), WiGR (Hu et al., 2021a), DAFI (Li et al., 2021c), Chen et al. (Chen and Chang, 2022), Ding et al. (Ding et al., 2021b), WiONE (Gu et al., 2021), Shi et al. (Shi et al., 2022), Caution (Wang et al., 2022e),
LT-WiOB (Zhou et al., 2022d), Wang et al. (Wang et al., 2020), Zhou et al. (Zhao et al., 2024c), CSI-GLSTN (Zhang et al., 2024e), HAR-MN-EFF (Shi et al., 2020c), DA-HAR (Sheng et al., 2023), CFLH (Wang et al., 2024f),
DualConFi (Xu et al., 2022), CLAR-Xiao et al. (Xiao et al., 2024a), CrossSense (Zhang et al., 2018), Sugimoto et al. (Sugimoto et al., 2023), CSI-MTGN (Zhang et al., 2023a), Wi-Fringe (Islam and Nirjon, 2020),
@@
WiFi2Radar (Nirmal et al., 2024), i-Sample (Zhou et al., 2024) , leaf,fill=yellow!15, text width=40em ] ] [ Data augmentation
2.3.5),fill=yellow!10 [ Wi-Dist (Zhang et al., 2024d), WiARGAN (Huang et al., [n. d.]), WiSGP (Liu et al., 2023b), Wang et al. (Wang et al., 2020), WGAN (Arjovsky et al., 2017), LESS (Zhang et al., 2023g), WiAG (Virmani and Shahzad, 2017), AirFi (Wang et al., 2022d),
CrossGR (Li et al., 2021b), Fido (Chen et al., 2020), Fidora (Chen et al., 2022), RFBoost (Hou and Wu, 2024), CLAR-Xiao et al. (Xiao et al., 2024a), Wi-Cro (Mao et al., 2024), freeGait (Yan et al., 2024c), SFTSeg (Xiao et al., 2022),
@@
WiTeacher (Xiao et al., 2023), adversarial autoencoder, (AAE) (Makhzani et al., 2015), freeLoc (Yan et al., 2024a) , leaf,fill=yellow!15, text width=40em ] ] [ Pseudo-labeling and
pure feature learning
2.3.6),fill=yellow!10 [ LAGER (Chen et al., 2024a), P3ID (He et al., 2024), OpenFi (Liu et al., 2024a), CLAR (Zhou et al., 2022a), TOSS (Zhou et al., 2022b), MAML (Finn et al., 2017), Zhang et.al (Zhang et al., 2023j), DA-HAR (Sheng et al., 2023), PIAS (Xiao et al., 2024b),
PAC-CSI (Su et al., 2023), 3D-ID (Ren and Yang, 2022), Abuhoureyah et al. (Abuhoureyah et al., 2023), Ma et al. (Ma et al., 2021), (Sheng et al., 2020) , leaf,fill=yellow!15, text width=40em ] ] ] [ Deployment
2.4), fill=teal!3 [ Transfer Learning
@@
2.4.1), fill=teal!7 [ XRF55 (Wang et al., 2024b), BullyDedect (Lan et al., 2024), WiLISensing (Ding et al., 2020b), DADA-AD (Chen et al., 2021), Brinke & Meratnia (Brinke and Meratnia, 2019b), WiTransfer (Fang et al., 2020),
@@
FewSense (Yin et al., 2022), ResMon (Zheng et al., 2023), RoGER (Bi et al., 2024), Hou et al. (Hou et al., 2022a), DASECount (Hou et al., 2022b) , leaf, fill=teal!10,, text width=40em ] ] [ Federated Learning
@@
2.4.2), fill=teal!7 [ WiFederated (Hernandez and Bulut, 2021), pFedBKD (Geng et al., 2023), CARING (Li et al., 2023), Qi et al. (Qi et al., 2023), Co-WiSensing (Zhang et al., 2023d), Zhang et al. (Zhang et al., 2021b),
@@
AdaWiFi (Zheng et al., 2024), Zhang et al. (Zhang et al., 2023f) , leaf, fill=teal!10,, text width=40em ] ] [ Continuous Sensing
@@
2.4.3), fill=teal!7 [ FewCS (Yang et al., 2023b), WiDrive (Bai et al., 2019), M-WiFi (Soltanaghaei et al., 2020), RISE (Zhai et al., 2021), CCS (Fu et al., 2024), CSI-ARIL (Zhang et al., 2023c) , leaf, fill=teal!10, text width=40em ] ] ] ]
Figure 1. Categories of Cross-Domain Wi-Fi Human Sensing.
2. Taxonomy

### v2

@@
Abstract
Wi-Fi sensing has emerged as a powerful non-intrusive technology for recognizing human activities, monitoring vital signs, and enabling context-aware applications using commercial wireless devices. However, the performance of Wi-Fi sensing often degrades when applied to new users, devices, or environments due to significant domain shifts. To address this challenge, researchers have proposed a wide range of generalization techniques aimed at enhancing the robustness and adaptability of Wi-Fi sensing systems. In this survey, we provide a comprehensive and structured review of over 200 papers published since 2015, categorizing them according to the Wi-Fi sensing pipeline: experimental setup, signal preprocessing, feature learning, and model deployment. We analyze key techniques, including signal preprocessing, domain adaptation, meta-learning, metric learning, data augmentation, cross-modal alignment, federated learning, and continual learning. Furthermore, we summarize publicly available datasets across various tasks—such as activity recognition, user identification, indoor localization, and pose estimation—and provide insights into their domain diversity. We also discuss emerging trends and future directions, including large-scale pretraining, integration with multimodal foundation models, and continual deployment. To foster community collaboration, we introduce the Sensing Dataset Platform (SDP) for sharing datasets and models. This survey aims to serve as a valuable reference and practical guide for researchers and practitioners dedicated to improving the generalizability of Wi-Fi sensing systems.
Index Terms:
@@ Abstract
I Introduction
@@
Wi-Fi sensing has emerged as a transformative technology that leverages ubiquitous Wi-Fi infrastructures for tasks beyond communication. By analyzing the variations in Wi-Fi signals such as Channel State Information (CSI) and Received Signal Strength Indicator (RSSI), researchers have demonstrated its potential in a wide range of applications, including activity recognition  [1, 2, 3, 4, 5, 6, 7], gesture recognition [8, 9, 10, 11, 12, 13, 14, 15, 16], indoor localization [17, 18, 19, 20, 21, 22, 23, 24], user authentication and identification [25, 26, 27, 28, 29, 30, 31], health monitoring [32, 33, 34, 35, 36, 37], fallen detection [38, 39, 40, 41, 42], pose estimation [43, 44, 45, 46, 47, 48, 49, 50], crowd counting [51, 52, 53, 54, 55], and more. Many survey papers have summarized this thriving research field, reviewing existing studies from perspectives such as methods and tasks [56, 57, 58, 59, 60, 61, 62].
Figure 1: Wi-Fi signals emitted from a transmitter propagate through both static environmental structures and dynamic human bodies before reaching the receiver. Human presence and movements alter the signal propagation paths, leading to measurable changes in the received signal. These variations can be leveraged for Wi-Fi sensing to interpret human movements.
While Wi-Fi sensing has shown great promise across a variety of applications, its effectiveness fundamentally relies on how wireless signals interact with the surrounding environment and the human body. Fig. 1 illustrates how human presence and motion influence the propagation of Wi-Fi signals. As shown in the figure, Wi-Fi devices are typically equipped with multiple transmit and receive antennas. Transmit antennas emit radio signals that propagate through the environment and reach the receive antennas via different paths, including line-of-sight, reflections from walls and furniture, and reflections from the human body. The presence and motion of a human subject alter the signal correlations on time, frequency or space,, resulting in changes to key signal properties such as Received Signal Strength (RSS), Angle of Arrival (AoA), and Time of Flight (ToF). By analyzing these signal changes, it is possible to infer meaningful information about human activities. Both handcrafted feature-based methods and data-driven deep learning models have been developed to map variations in Wi-Fi signals to human presence, location, motion, respiration, and more—thus enabling a wide range of Wi-Fi sensing capabilities.
(a) Device heterogeneity
(b) Human body shape diversity
(c) Environment diversity
Figure 2: Wi-Fi sensing generalization is primarily hindered by three key factors, i.e., device heterogeneity, human body diversity, and environmental diversity.
Figure 3: Growth of research in Wi-Fi generalization: from a handful of studies between year of 2015 and 2018 to a surge of publications since 2019.
However, one critical challenge remains unresolved: generalization—the performance of Wi-Fi sensing often degrades when applied to new users, devices, or environments due to significant domain shifts. This performance gap between research prototypes and practical systems arises from the inherent variability in Wi-Fi sensing conditions, which introduces considerable domain shifts across deployment scenarios. As illustrated in Fig. 2, generalization is primarily hindered by three key factors, including device heterogeneity, human body diversity, and environmental diversity. First, device heterogeneity introduces substantial signal variation. Wi-Fi signals originate from a wide range of devices—including routers, laptops, and mobile phones—that differ in protocol versions, operating channels, bandwidth, antenna design, physical orientation, and chipset configurations. These variations affect the measured CSI and RSS, often rendering models trained on one device ineffective on another. Second, human body diversity introduces non-negligible discrepancies in signal interactions. Variations in height, weight, body composition, and even clothing materials can alter the propagation and absorption of Wi-Fi signals. As a result, two individuals performing the same action at the same location may produce significantly different signal patterns, posing challenges to consistent activity recognition. Third, environmental diversity in indoor scenes further exacerbates generalization difficulties. Differences in room layouts, wall materials, furniture placement, and device positions reshape the signal propagation paths and multipath profiles, leading to domain-specific variations in signal features such as RSS, AoA, and ToF. These changes severely limit the Wi-Fi sensing systems from generalizing beyond their training conditions.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6em,align=left,fit=band,font=,, where level=2text width=8.2em,align=left,fit=band,font=,, where level=3text width=9em,font=,, where level=4text width=8em, font=,, [ Wi-Fi Human Sensing Generalization Categories, ver [ Experimental
@@
II-A) , fill=orange!5 [ Distributed Antennas
II-A1), fill=orange!10,text width=9em [ WiNDR [63], WiCross [64], WiSDAR [65] , leaf,fill=orange!10, text width=42em ] ] [ Distributed Devices
II-A2),fill=orange!10,text width=9em [ ReWiS [66], Widar 3.0 [67], OneFi [68], WiTraj [69] , leaf,fill=orange!10, text width=42em ] ] [ Scaling Up Dataset
II-A3),fill=orange!10,,text width=9em [ XRF55 [70], Widar 3.0 [67], Zhang et al.[71], MM-Fi[72], CSI-Bench [73] , leaf,fill=orange!10, text width=42em ] ] ] [ Signal
@@
II-B), fill=violet!5 [ Signal Analysis
Methods (§II-B1), fill=violet!10,text width=9em [ WiHand [74], CLAR [75], Gait-Enhance[76], Zhuo et al.[77], Abuhoureyah et al. [78], leaf, fill=violet!15, text width=42em ] ] [ Advanced Handcrafted
Indicators (§II-B2), fill=violet!10, text width=9em [ WiHGR [79], PhaseAnti [80], DPSense [81], HandGest [82], Yu et.al. [83], AF-ACT [84], RoSeFi [85],
Wang et.al. [86], WiGesFree[87], LiKey [88], Bu et.al. [89], Zhang et.al. [90] , leaf, fill=violet!15, text width=42em ] ] [ Motion Indicators
II-B3), fill=violet!10,text width=9em [ Sharp [91], OneSense [92], Yin et al.[93], Wi-learner [94], Niu et al. [95], and LAGER [96], WiLife [97],
Mini-Batch Alignment [98, 99], WiTraj [69], Widar 3.0 [67]. Zhang et al. [100], WiTransformer [101],
Shi et al. [102], Bulugu [103]. GaitID [30], OneFi [68], XFall [104], WiGesture [105], Shi et al. [106],
Wang et al. [107], Widir [108], Wang et al. [109], WiNDR[63], WiDIGR [110], Wi-PIGR [111],
WiDetect [112], Zhu et al. [113], leaf, fill=violet!15, text width=42em ] ] [ Angle Indicators
II-B4), fill=violet!10,text width=9em [ Li et al. [114], EasyTrack [115], 3D-ID [116], leaf, fill=violet!15, text width=42em ] ] ] [ Feature
@@
II-C),fill=yellow!5 [ Domain Alignment
@@
II-C1),fill=yellow!10,text width=9em [ WiHARAN [117], DA-HAR [118], Zhang et al. [119], i-Sample [120], JADA [121], CrossGR [122], Wi-Adaptor [123],
EI [124], Khattak [125], WiCAR [126], GESFI [127], ADA [128], DATTA [129], WiSR [130], Yin et al.[93],
INDG-Fi [131], WiCross [64], Berlo et al. [132], DINN [133], freeGait [134], DAFI [135], Shi et al. [106], Fidora [136],
MMD [137], AdaPose [138], BSWCLoc [139], Kang et al. [140], WiAi-ID [141], Wi-CHARS [142], Mehryar [143],
AdapLoc [144], PIAS [145], CDD [146], WiCAU [147], WiSDA [148], FDAS [149], CDFi [150], LAGER [96],
Zhan & Wu [151], AirFi [152], WiLISensing [153], MetaGanFi [154], Zhang et al. [155] , Wi-Cro [156], WiTeacher [157],
CrossSense [158], Sugimoto et al. [159], CSI-MTGN [160], Wi-Fringe [161], XRF55 [70], Wi-Chat [162], WiFi2Radar [163] , leaf,fill=yellow!15, text width=42em ] ] [ Component Disentangle
II-C2),fill=yellow!10, text width=9em [ Elujide et al. [164], UH-Sense [165], Person-in-WiFi [45], Wi-Piga [166], CrossID [167], LT-WiOB [168], CATFSID [169],
TransferSense [170], UniFi [171], WiOpen [172], NCA [173], SFTSeg [174], DualConFi[175], CFLH [176], Xiao et al. [177],
3D-ID [116], MaP-SGAN [178], CrossFi [179], Wang et al. [180], WiGR [181], LESS [182], WiGesID [183] , leaf,fill=yellow!15, text width=42em ] ] [ Metric Learning for
Zero/Few-shot Learning
II-C3),fill=yellow!10, text width=9em [ LT-WiOB [168], CATFSID [169], TransferSense [170], UniFi [171], WiOpen [172], NCA [173], SFTSeg [174],
DualConFi[175], CFLH [176], Xiao et al. [177], 3D-ID [116], MaP-SGAN [178], CrossFi [179], Wang et al. [180],
WiGR [181], LESS [182], WiGesID [183], AFEE-MatNet [184], HAR-MN-EF [185], MatNet-eCSI [186], Zhou et al. [187],
WiLiMetaSensing [188], ReWiS [66], Ding et al. [189], WiONE [190], CAUTION [191], FewCS [192], Zhang et al. [193] , leaf,fill=yellow!15, text width=42em ] ] [ Meta Learning for
Few-shot Learning
II-C4),fill=yellow!10,text width=9em [ MAML [194], Huang et al. [195], TOSS [196], MetaLoc [197], MetaFormer [198], OneSense [92], Wi-learner [94],
@@
RF-Net [199], WiGR [200], MetaGanFi [154], Zhang et al. [201] , leaf,fill=yellow!15, text width=42em ] ] [ Data Augmentation
and Data Synthesize
II-C5),fill=yellow!10,text width=9em [ CrossGR [122], Wang et al. [180], SIDA [202], WiARGAN [203], AFSL-HAR [204], LESS [182], Wi-Dist [205],
CsiGAN [206], Fidora [136], Fido [207], freeLoc [208], freeGait [134], Xiao et al. [177], Zhou et al. [209], CSI-Net [210],
BullyDetect [211], P3ID [212], OpenFi [213], Cutout [214], AirFi [152], RFBoost [215], WiSGP [216], WiAG [217],
@@
Wi-Cro [156], SFTSeg [174], OneFi [68], OneSense [92], Wang et al. [218] , leaf,fill=yellow!15, text width=42em ] ] [ Pseudo Labeling
II-C6),fill=yellow!10,text width=9em [ CLAR [75], LAGER [96], Zhang et al. [219], DA-HAR [118], Zhang et al. [100] , leaf,fill=yellow!15, text width=42em ] ] ] [ Model
Deployment
II-D), fill=teal!3 [ Transfer Learning
@@
II-D1), fill=teal!7,text width=9em [ XRF55 [70], BullyDedect [211], WiLISensing [153], DADA-AD [220], Brinke & Meratnia [221], WiTransfer [222],
@@
FewSense [223], ResMon [224], RoGER [225], Hou et al. [226], DASECount [227] , leaf, fill=teal!10,, text width=42em ] ] [ Federated Learning
@@
II-D2), fill=teal!7,text width=9em [ WiFederated [228], pFedBKD [229], CARING [230], Qi et al. [231], Co-WiSensing [232], Zhang et al. [119],
@@
AdaWiFi [233], Zhang et al. [234] , leaf, fill=teal!10,, text width=42em ] ] [ Continual Learning
@@
II-D3), fill=teal!7,text width=9em [ FewCS [192], WiDrive [235], M-WiFi [236], RISE [237], CCS [238], CAREC [239], CSI-ARIL [240] , leaf, fill=teal!10, text width=42em ] ] ] ]
Figure 4: We systematically review and categorize nearly a decade of research on Wi-Fi sensing generalization, covering over 200 publications since 2015. Distinct from previous surveys, we organize the literature along a four-stage sensing pipeline—device deployment, signal preprocessing, feature learning, and model deployment.
To enhance the generalization of Wi-Fi sensing systems, a diverse array of techniques has been proposed. As shown in Fig. 3, research in this field has expanded significantly, evolving from a few studies published between 2015 and 2018 to a surge of recent work focused on improving Wi-Fi generalization. These efforts can be broadly classified into four key stages of Wi-Fi sensing systems: experimental setup, signal preprocessing, feature learning, and model deployment.
  • Experimental Setup: The primary function is preparation and placement of devices and antennas. There are some approaches that improve generalization by strategically placing antennas around the target area or the human subject [63, 64]. This spatial diversity helps mitigate the impact of orientation changes and body blocking effects on signal reception. For instance, placing transceivers at multiple angles can reduce performance degradation caused by changes in a subject’s facing direction.
  • Signal Preprocessing: The primary function is extraction of sensing information from Wi-Fi signals, such as denoising and signal transformation. Researchers have explored the use of domain-invariant features that are closely linked to human motion. Examples include Doppler shift and body velocity profile [67], which are more resilient to environmental variation than raw CSI or RSS data.
  • Feature Learning: The primary function is training Wi-Fi sensing models using data. A rich set of machine learning techniques has been employed to bridge domain gaps. These include domain adaptation [241], domain generalization [242], meta-learning [194], and synthetic data generation, all aiming to learn features that are transferable across devices, subjects, and environments.
  • Model Deployment: The primary function is deploying trained models to real-world environments. Techniques such as model fine-tuning, federated learning, and continual learning have been used to adapt pre-trained models to new settings with minimal labeled data.
(a) Distributed Antennas
(b) Distributed Devices
(c) Device Placement
(d) Scaling up datase
Figure 5: In the device deployment stage, Wi-Fi generalization can be enhanced by distributing antennas to mitigate the impact of user orientation, deploying devices more widely or optimizing their placement to increase coverage, and collecting more diverse datasets to support the training of more robust and generalizable models.
This survey systematically reviews and categorizes nearly a decade of research on Wi-Fi sensing generalization, covering over 200 publications since 2015, as shown in Fig. 4. Distinct from previous surveys [243, 244], we organize the literature along a four-stage sensing pipeline—experimental setup, signal preprocessing, feature learning, and model deployment—to provide a structured and comprehensive view of existing techniques. This stage-wise organization enables researchers and practitioners to quickly identify methods suited to their specific challenges. Furthermore, we curate and summarize a wide array of publicly available datasets, offering guidance for researchers seeking appropriate benchmarks to train or evaluate generalizable models. To support future research, we also explore emerging trends, including multimodal sensing, large pre-trained models, and data-efficient learning paradigms, that are poised to shape the next generation of Wi-Fi sensing technologies. We hope this survey will serve as a comprehensive and practical resource for both newcomers and experienced researchers seeking to understand, implement, or advance generalization techniques in Wi-Fi sensing.
The remainder of this paper is organized as follows. Section II provides a stage-wise taxonomy of generalization techniques in Wi-Fi sensing, covering key methods and representative studies across device deployment, data preprocessing, feature learning, and model deployment, respectively. Section III presents a comprehensive overview of publicly available Wi-Fi sensing datasets, categorized by sensing tasks, including action recognition, pose estimation, crowd counting, etc. Section IV discusses promising future research directions, including multimodal sensing, large-scale model pretraining, and continual learning. Finally, Section V concludes the paper by summarizing key insights on the path forward for building more generalizable Wi-Fi sensing systems.
II Taxonomy